Please List Steps, show all work & explain logic

24. Calculate [Ag at 298K for the following reaction as written using the conditions listed below: AgCl (s) Ag (aq) +Cl (aq) [Cl] 1.0 x 10 M AG 58.496 kJ/mol AGO +55.600 kJ/mol A. 1.0 x 10 10 M B. 1.0 M C. 1.3 x 100 M D. 7.9 x 10 M E. There is no Ag (aq) in the solution.

Calculate [Ag^+] at 298K for the following reaction as written using the conditions listed below:  AgCl (s) rightarrow Ag^+ (aq) + Cl^-(aq)  [Cl^-] = 1.0 times 10^-10 M  delta G = -58.496 kJ/mol  delta G degree = +55.600 kJ/mol  1.0 times 10^-10 M  1.0 M  1.3 times 10^-5 M  7.9 times 10^-8 M  There is no Ag^+ (aq) in the solution.
Answer #1

delta G = delta Go + R*T*ln Q

delta G = delta Go + R*T*ln ([Ag+][Cl-])

-58496 J = 55600 J + 8.314* 298 * ln ([Ag+]*1.0*10^-10)

ln ([Ag+]*1.0*10^-10) = -46.0515

[Ag+]*1.0*10^-10 = 1.000*10^-20

[Ag+] = 1.0*10^-10 M

Answer: A


answered by: ANURANJAN SARSAM
Answer #2
[REMOVED]